OpenStudy (anonymous):

What is the area of a circle with a diameter of twenty-two meters? Use twenty-two sevenths for pi. three hundred seventy and six-sevenths meters squared three hundred seventy-four meters squared three hundred seventy-seven and one-seventh meters squared three hundred eighty and two-sevenths meters squared

OpenStudy (anonymous):

@zpupster @Catseyeglint911 @cwrw238 @Lyssa123

OpenStudy (solomonzelman):

if diameter is 22 meters, then the radius is 11. A= π r²

OpenStudy (solomonzelman):

Use a 3.14 approximation for π

OpenStudy (anonymous):

ok multiply 11*11 and you get 121 and now you multiply 3.14 and get 379.94

OpenStudy (anonymous):

is it correct

OpenStudy (solomonzelman):

Yes, totally !

OpenStudy (anonymous):

but its not in the question thats what got me confused

OpenStudy (anonymous):

help me @zpupster

OpenStudy (solomonzelman):

Ohh, you have to use 22/7 for π. So then the radius is still 11, and so A = π r² = (22/7) × 11² = (22/7) × 121 = 121 × (22/7) = (121 × 22)/7 = 2662 / 7 = 380 and 6/7 .